OpenStudy (anonymous):

If u is nonzero and u, v and w are vectors, is it valid to cancel u from both sides of the equation u x v = u x w and conclude that v=w? Explain your reasoning.

OpenStudy (anonymous):

no it isnt in this case u can say u x v = u x w hence u x v - u x w = 0 u x (v-w) = 0 so we see that u is collinear to (v-w) hence u = t(v-w) where t is any real parameter

OpenStudy (anonymous):

get it?

OpenStudy (anonymous):

What's collinear?

OpenStudy (anonymous):

means the two vectors run parallel to each other

OpenStudy (anonymous):

oh okay make sense thanks

OpenStudy (anonymous):

actually one more question how do you know that u x (v-w) is parallel? it's cross product

OpenStudy (anonymous):

when the cross product of 2 vectors is 0 that mean theyre parallel to each other see a x b = absint where t is the angle b/w the vectors a x b = 0 and vectors are both nonzero so => sin t = 0 the angle bw the vectors is zero

OpenStudy (anonymous):

oh yeah okay thanks a lot man